OpenStudy (anonymous):

Numbers p and q are negative integers. Which statements are always true? Choose all answers that are correct. A. p + q is a negative integer B. p • q is a positive integer C. p – q is a negative integer D. is a negative integer

OpenStudy (zkoqmosii):

B

OpenStudy (zkoqmosii):

Because negative * negative = positive. None of the rest are right.

OpenStudy (freckles):

what about a neg+neg...?

OpenStudy (freckles):

I do agree neg*neg is pos but neg+neg is going to give you a more negative number

OpenStudy (zkoqmosii):

Yeah you're right.

OpenStudy (freckles):

assume a and b are positive -a+-b=-(a+b) a+b is positive since a and b are positive but there is a negative in front of (a+b) which makes the final result negative

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

so the answer would be A and..... what? PS. D is this...

OpenStudy (freckles):

a negative times a negative gives you a what @goopyfish908

OpenStudy (anonymous):



OpenStudy (anonymous):

a negative?

OpenStudy (freckles):

so -2 times -1 would give you -2 ? or 2?

OpenStudy (freckles):

-2*(-1) should be 2 not -2

OpenStudy (anonymous):

oh, ok

OpenStudy (freckles):

so a neg*neg gives you a ?

OpenStudy (anonymous):

posotive?

OpenStudy (freckles):

yep

OpenStudy (freckles):

neg+neg=neg and neg*neg=pos

OpenStudy (anonymous):

ok, so... the answer is A and B?

OpenStudy (anonymous):

or, ABC

OpenStudy (anonymous):

????

OpenStudy (anonymous):

IM confused

OpenStudy (anonymous):

IM confused

OpenStudy (freckles):

just the one where it says neg+neg=neg and the one where it says neg*neg=pos

OpenStudy (anonymous):

sooooo............

OpenStudy (freckles):

the neg-neg=neg is false because for example -2-(-2)=-2+2=0 and 0 is not neg neg/neg=neg is also false example -4/-2 is 2 and 2 is positive not negative